OpenStudy (anonymous):

What is the slope of a line that passes through the point (−2, 3) and is parallel to a line that passes through (3, 7) and (−2, −8)? One solution of the equation 2x 2 − x − 10 = 0 is . What is the other solution? Given the function g(x) = −4x + 5, find g(−4) 3 questions I need help on if anyones willing to help.

OpenStudy (anonymous):

The other solution is 5/2, and Id prefer if anyone would walk me through it instead of giving me the answer.

OpenStudy (anonymous):

I got 15/5

OpenStudy (anonymous):

The negatives should cancel out since both the numerator and denominator are negative

OpenStudy (anonymous):

For which one?

OpenStudy (anonymous):

Slope=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

For the lines to be parallel, both equations have the same slope, thus they never touch

OpenStudy (anonymous):

^For Question #1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I really dont understand

OpenStudy (anonymous):

Which part?

OpenStudy (anonymous):

so im finding the the parallel line instead of the slope?

OpenStudy (anonymous):

well, you're finding the slope of the line, which happens to be parallel to another line

OpenStudy (anonymous):

So to find the slope I find the parallel first and then what?

OpenStudy (anonymous):

Do you know how to find the slope with two points?

OpenStudy (anonymous):

no. Thats why I asked whoever would help me to walk me through it

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

slope w/ two points can be found doing (y2-y1)/(x2-x1) In this case x1=3, y1=7, like point (3, 7) and x2 = -2 and y2= -8 like (−2, −8)

OpenStudy (anonymous):

That gives you (-8-7) *y2-y1* divided by (-2-3) *x2-x1*

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

ok so then its -15/-5= 3. Right?

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

oh ok!

OpenStudy (anonymous):

lol do you get the parallel line & slope relation?

OpenStudy (anonymous):

^ Parallel lines have the same slope

OpenStudy (anonymous):

yes I do ty. And on the others?

OpenStudy (anonymous):

K

OpenStudy (anonymous):

for number 3, when you see g(-4), it means you substitute (-4) in for any x values in the equation

OpenStudy (anonymous):

oh! ok! I thought it ment to replace the g

OpenStudy (anonymous):

yeah, it's a little confusing lol

OpenStudy (anonymous):

If you put in the values, you should get 21 for #3

OpenStudy (anonymous):

I got that but what do I do with the -4 by the g? (or does that dissapear?)

OpenStudy (anonymous):

Technically, it disappears because it means only when x=-4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but you could leave g(-4)

OpenStudy (anonymous):

ok. and the last one?

OpenStudy (anonymous):

what's one of the solutions for #2? The question kind of ends mid-sentence

OpenStudy (anonymous):

the solution thats given is 5/2.

OpenStudy (anonymous):

hmm one sec

OpenStudy (anonymous):

This one's hard -.-'

OpenStudy (anonymous):

I got it!

OpenStudy (anonymous):

So, since the equation is a polynomial, you can factor the equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you know how to factor or do you want a walkthrough for it?

OpenStudy (anonymous):

@xNoahx?

OpenStudy (anonymous):

um Im not 100% sure

OpenStudy (anonymous):

on factoring that is

OpenStudy (anonymous):

all right

OpenStudy (anonymous):

in factoring, you want to make two binomials that, when multiplied, equal to 2x 2 − x − 10 = 0

OpenStudy (anonymous):

how would you do that?

OpenStudy (anonymous):

To start, you know that one set has to be (2x+ or -) and (x+ or -)

OpenStudy (anonymous):

ok so am I changing the x? Or am I finding a number to add/subtract the other numbers by?

OpenStudy (anonymous):

do you know the FOIL method (first outer inner last) of multiplying?

OpenStudy (anonymous):

yes I do. but how would I use that in this equation?

OpenStudy (anonymous):

okay, to get the leading 2x^2, you multiply the first term of the two binomials

OpenStudy (anonymous):

the only way to get 2x^2 would be by having 2x in one binomial and x in the second

OpenStudy (anonymous):

ok so It would be (2x+-#) (x+-#)?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

but where do we get the number from?

OpenStudy (anonymous):

to find the # numbers, you have to find two number that, when multiplied by eachother, they equal 10

OpenStudy (anonymous):

so 2 and 5?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

that gives you (2x-5) and (x-2)

OpenStudy (anonymous):

ok so i would be (2x+-2) (x+-5)?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

I got 2x2-4x-5x-10

OpenStudy (anonymous):

so, each equation, by default, is set to 0, so it's 2x-5=0 and x-2=0

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

you said foil so I did that. was that wrong?

OpenStudy (anonymous):

no you're right

OpenStudy (anonymous):

I might have made a mistake. I'll check again

OpenStudy (anonymous):

ok so I add together -4x and -5x right?

OpenStudy (anonymous):

okay, the two binomials are x-2 and 2x+ 5 (not -5)

OpenStudy (anonymous):

oh oops

OpenStudy (anonymous):

lol no my fault

OpenStudy (anonymous):

so its 2x2-4x+5x-10?

OpenStudy (anonymous):

Shoot

OpenStudy (anonymous):

-.-' I messed up again. It's really (2x-5) and (x+2).

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

sorry about that :(

OpenStudy (anonymous):

so its 2x2+4x-5x-10?

OpenStudy (anonymous):

? It should be 2x^2+4x-5x-10

OpenStudy (anonymous):

thats what I put XD

OpenStudy (anonymous):

-.- I'm losing it, aren't I?

OpenStudy (anonymous):

XD its ok! so I add together 4x and -5x?

OpenStudy (anonymous):

yeah, it simplifies to -x, like the equation

OpenStudy (anonymous):

like before, you set each equation to 0, so it's 2x-5=0 and x-2=0

OpenStudy (anonymous):

then, you solve for x

OpenStudy (anonymous):

so its 2x2-x-10=0?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

is it x=2x2-10?

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

2x-5=0 and x-2=0 give you the answers x= 5/2 and x= 2

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

when you get the 2 binomials of the equations, you use those to find x's solutions

OpenStudy (anonymous):

ok ty!

OpenStudy (anonymous):

No problem! I'm glad you get it ^^